2(Cos 2 73 o +cos 2 47 o )-154 o it seems very hard please help me

2(Cos2 73o +cos247o)-154o    
it seems very hard please help me

Grade:11

1 Answers

somi teez
6 years ago
Hi,
i will help..take ease....please go through each step
so,
let A=73 degree and B=47 degree  
 we find that , 154=180-(73-47) in terms of degree
 Expression is now,
            =2( cos^2 A + cos^2 B ) - cos [ pi - (A - B) ]
           = 2cos^2A +2cos^2B + cos(A - B)                                              [ using cos (pi-X) = - cosX ]
So,
           =2cos^2A +2cos^2B + cos(A - B)
           = (2 cos^2A - 1) + (2 cos^2B - 1) + cos(A - B) +2 
           = cos 2A + cos 2B + cos(A - B) +2 
           = cos146 + cos94 + cos26 +2 
Now,
          =[cos 146 + cos 94 ] = 2cos120 * cos26                                      { cosC+cosD = 2cos(C+D)/2*cos(C-D)/2 }
So now expression,
          = 2cos120 * cos26 + cos26 + 2 
          = 2(-1/2)*cos26+cos26+2                                                            we know  cos 120 = - cos 60 = -1/2 

          = -cos26+cos26+2 = 2.   
i hope you understand...just go with the formulas
approve it if you agree 
BEST OF LUCK!!!!